Kingsport containers company makes a single product that is subject to wide rates....

Steps if possible

@ https//newconnect.mheducation.com/flow/connect.html Ch 2 Homework 5 Exercise 2-11 Varying Plantwide Predetermined Overhead Rates [LO2-1, LO2-2, LO2-3] 17 points Kingsport Containers Company makes a single product that is subject to wide seasonal varniations in demand. The company uses a job-order costing system and computes plantwide predetermined overhead rates on a quarterly basis using the number of units to be produced as the allocation base. Its estimated costs, by quarter, for the coming year are given below eBook Print Reterences FirstSecond ThirdFour $240,000 $120,000 % 60,000 $180,000 Direct materials Direct labor 120,000 60,000 30.00 e,00 Manufacturing overhead Total manufacturing costs (a) 529 005386,000 284,909 120,000 60,0e9 ,000 9e,000 6.439.4 Number of units to be produced (b) Estinated unit product cost (a) (b) 4.92 Management finds the variation in quarterly unit product costs to be confusing and difficult to work with. it has been suggested that the problem lies with manufacturing overhead because it is the largest element of total manufacturing cost Accordingly, you have been asked to find a more appropriate way of assigning manufacturing overhead cost to units of product Required: 1. Assuming the estimated vanable manufacturing overhead cost per unit is sO.40, what must be the estimated total fixed manufacturing overhead cost per quarter? 2 Assuming the assumptions about cost behavior from the first three quarters hold for the fourth quarter? 3. What is causing the estimated unit product cost to fluctuate from one quarter to the next? 4. Assuming the company computes one predetermined overhead rate for the than calculate the unit product cost for all units produced during the year constant, what is the estimated unit product cost ea ather than coui uartery overhead rates computing quarterly overhead rates, calculate the unit product cos foral ntsproduced ciuing the e Complete this question by entering your answers in the tabs below c Prev 5 of 6 Score ansiwer >
Answer #1

1.we shall estimate the variable manufacturing overhead cost per unit through high low method:

as per high low method = variable portion of cost = (highest cost - lowest cost ) / (highest units - lowest units)

=> (230,000 - 194,000) / (120,000 - 30,000)

=>36,000 / 90,000

=>$0.40. per units.

now,

consider the highest level 120,000 units with a total overhead cost of $230,000.

=> variable overhead + fixed overhead = 230,000

=>(120,000*0.40) + fixed overhead = 230,000

=>fixed overhead =>fixed overhead = 230,000 - 48,000

=>182,000.

estimated fixed manufacturing overhead cost per quarter = $182,000

2.unit product cost for fourth quarter:

direct materials 180,000
direct labour 90,000
manufacturing overhead ($0.40*90,000units)+182,000 218,000
total manufacturing cost $488,000
number of units 90,000
estimated unit product cost (488,000/90,000) $5.42

3.Fixed manufacturing overhead.

SInce the fixed manufacturing overhead is being estimated on a quarterly basis, based on the number of units to be manufactured, there is a change in unit product cost from quarter to quarter.

4.total units produced in the year = 120,000+60,000+30,000+90,000 =>300,000

total fixed overhead = (182,000 * 4 quarters) =>728,000

total variable overhead = ($0.40*300,000)=>120,000.

total manufacturing overhead =(728,000+120,000)=>$848,000

manufacturing overhead cost per unit = ($848,000/300,000 units) =>$2.83.